OpenStudy (anonymous):

PLEASE help!! Medal & Fan! f y = 9x - 7, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, - 7) , (1, 2) , (-1, -16)} {(-7, 0), (2, 1), (-16, -1)} {(1, 9), (2, 7), (3, 16)} {(7, 9), (8, 10), (9, 11)}

OpenStudy (anonymous):

it would be the first one because -7= 9*0 - 7 is 0 2 = 9*1 -7 IS 9 -16 = 9*-1 -7 is -9